OpenStudy (anonymous):

Which of the following are solutions to the equation below? (2x + 3)2 = 11

OpenStudy (anonymous):

all you have to is distribute and solve, then you will have your solution

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes... would it be 2x+9=11

OpenStudy (anonymous):

Hmm are you sure? 2(2x)+2(3)=11

OpenStudy (anonymous):

4x+6=11?

OpenStudy (anonymous):

yes great!

OpenStudy (anonymous):

Now solve 4x+6=11. We have to get the number with the variable by itself

OpenStudy (anonymous):

Ill finish! 4x+6=11 -6 -6 ____________________ 4x=5 x=?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

close 4/5

OpenStudy (anonymous):

great effort though :)

OpenStudy (anonymous):

thanks